xy=21
y=3x-2
could you please help me with this non-linear system of equations?
I started the problem x(3x-2)=21 then I got 3x^2-2x-21=0, but I didn't know where to go from there.
Thank you for your help.

Answer by checkley71(8403)

3X~2-2X-21=0 OR (3X+7)(X-3)=0 THEREFORE X=-7/3 AND X=3